OpenStudy (anonymous):

Solve the equation. helps help medal will be given! A. 4= Square root p-2 B. spuare root q-9=7

OpenStudy (anonymous):

\[4=\sqrt{p-2}\]

OpenStudy (anonymous):

p-2=4^2 p-2=16 p=18

OpenStudy (anonymous):

square both side

OpenStudy (anonymous):

q-9=49 q=58

OpenStudy (anonymous):

16=p-2 16+2=p p=18

OpenStudy (anonymous):

\[\sqrt{q-9=7}\]

OpenStudy (anonymous):

square both side q-9=49 q=49+9=58

OpenStudy (anonymous):

so A isnt right i cant pick that one

OpenStudy (anonymous):

A and B both are right